OpenStudy (anonymous):

Need to know how to find a missing number when dividing a fraction ie 16/5 divied by ? = 16/2

OpenStudy (amistre64):

(16/5)/x = 16/2 ; multiply each side bythe reciprocal of (16/5); which is just (5/16) 1/x = (16/2)(5/16) 1/x = 5/2 ; flip each one over to get: x = 2/5